Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 584     Entry time: 05 Dec 2018, 11:41
Author: Adriano Violante 
Location: General 
Severity: Information 
Subject: Status report after dispatch meeting 09.30 

SPB

  • pulse arrival time measurements to be done this week.
  • second Gotthard detector small issues being fixed (online data viewer fixes - ongoing)

FXE

SQS

SCS

MID

  • Loop 10 partially needed. Smaract stages are working withing loop10 since yesterday

HED

XRO

XPD

VAC

AE

  • AE agreed to power down the PLCs in SASE1 experiments and optical laser before the power interruption. Dates are being rediscussed since radiation damages tests are foreseen for Monday and Tuesday! (10-11.12). TBC.
  • The synch of the Gotthard detectors with 1 trigger does not work for long trains (synch is lost). Multiple triggers are needed and also possible (AGaller is already using them). AE can give support also to the other users.

  EET

ITDM

DET

  • SCS fast CCD temperature drift issue being addressed.

CAS

  • Tests of the 2.3.0. Karabo version are ongoing.  Karabo update for SASE1 will be deployed after the  PLCs are back again (ca. 2 days after the power is back). Proposed date by CAS is 19.12. but needs to be discussed again once the power is back.

AoB

ELOG V3.1.4-7c3fd00